# Tensor networks for quantum computing

### v1

@@ IV Quantum circuit synthesis
TNs offer a promising general pathway for addressing these challenges [128]. TN architectures, such as MPS and MPO, feature an inherently geometrical layout provided by the prescribed decomposition/representation of general tensor formats. It turns out that for circuit descriptions, which inherently are TNs, such layouts are particularly attractive due to the ease in attaining connectivity within adjacent sets of qubits. Although this feature is exploited in studying long/short range interactions in condensed matter systems, it naturally provides a partial solution to the problem of circuit synthesis in terms of realizing native hardware connectivity. By representing quantum states and operations as TNs and further casting them as circuits, one can simplify both circuit design as well as address the problem of compiling arbitrary unitaries into natively realizable gates. This can reduce the complexity of quantum circuits and enable more efficient synthesis.
@@ IV.2 Examples of a TN state and operator preparation
Using the introduced prescription, we provide examples of important TN archectures in the literature, and demonstrate how they can be mapped to quantum circuits. Figure 3 depicts how one converts an MPS into a quantum circuit using the steps from the above subsection. The key element in this procedure is that an MPS can be exactly transformed into a canonical form, including only isometric tensors, which is a unique feature of an MPS—this is not the case for the higher-dimensional TNs like PEPS. Preparing an MPS as a quantum circuit was first introduced in [131], and has since been theoretically explored by many others [132, 133, 134, 135, 136]. An experimental realization of an MPS has been extensively explored across various quantum computer architectures [137, 138, 139, 140].
@@
The PEPS, however, generally lack a canonical exact isometric form due to the higher connectivity inherent in their underlying structure. In other words, one cannot exactly map a PEPS onto a quantum circuit without an exponential amount of postselection. However, one can create a subclass of PEPS, referred to as isoPEPS, to establish a proper connection to quantum circuits. In the isoPEPS class, each tensor, or set of tensors, respects the isometric condition, making it equivalent to a quantum circuit [88, 142, 143]. The question of undestanding the varational power of isoPEPS and the way how it is compared to PEPS and to 2dMERA and how they can be realized experimentally on a digital quantum computer is now an active area of research.
Among all tensor network methods for classical simulation of quantum systems, the MERA stands out as the most natural representation of a quantum circuit: it was originally envisioned in the reverse direction of a quantum circuit in which all isometries are embedded in unitaries. It features an intrinsic robustness to noise and does not suffer from barren plateaus [144, 145, 146]. The MERA has recently been demonstrated on the ion-trapped digital quantum computer to probe the criticality of many-body systems [147, 148, 149].
@@ IV.3 Implementation techniques
As mentioned in Sec. 4.1, TNs cannot be directionality compared to real quantum circuits. Because of this, we must choose how tensors are executed in time, which then is related to the quantum circuit architecture. This freedom allows TN states to be prepared holographically. Introduced first in [137], a holographic preparation of a TN state takes a spatial dimension of the physical state and prepares it sequentially in time, with the important feature that once a circuit is done preparing one site, the qubit can be reused to prepare the next site.
@@
Beyond the exact implementations of mapping a TN to a quantum circuit described above, various variational methods are employed to prepare an approximate TN state [157, 158, 159]. The basic idea is to numerically optimize a variational ansatz to approximate the target state, typically resulting in a shallower quantum circuit. In addition to preparing a TN state, the work [160] used variaional methods to instantiate a generic quantum circuit where the TN formulation is leveraged as an efficient back-end in the compilation workflows.
@@ V Quantum error correction and mitigation
Quantum error correction (QEC) plays a critical role in safeguarding quantum information from errors caused by decoherence, dissipation, and control inaccuracies [127]. In this section, we discuss two main topics: (i) the parallels between quantum error-correcting codes and TNs, along with (ii) the application of TNs in decoding, including the formal reduction of optimal decoding to a TN contraction [161].
@@ V.1 Tensor-network codes
The first major intersection between TNs and QEC was the establishment of a formal correspondence between certain QEC codes and the associated TNs [161]. Convolutional codes, concatenated block codes, and topological codes, for example, can be respectively represented by the MPS, tree TNs, and PEPS. This insight allows for importing well-established tools of TNs to represent and analyze QEC codes. In particular, the problem of an optimal decoding of an QEC code (i.e. diagnosing errors, see the discussion in Section V.2) was shown to be formally equivalent to contracting an associated TN [161] approach. For convolutional and contatenated block codes, which have efficiently contractable TNs, this correspondence thereby provides the efficient optimal decoders.
The QEC-TN correspondence was later generalized by the introduction of TN stabilizer codes [162] and the ‘quantum lego’ formalism [163, 164, 165] for constructing large codes out of a finite set of smaller ‘seed’ codes. The seed codes are representend by small tensors that are combined into a TN to represent a larger code. This construction was shown to be universal in the sense that any qudit QEC code can be represented by a TN built out of three elementary seed tensors [163]. As before, TN codes come naturally equipped with optimal decoders that are evaluated by contracting a TN [162].
As an important special case, TNs provide a powerful formalism for representing arbitrary topological and holographic codes [162, 166]. For example, Ref. [166] showed how simple modifications of the TN for a topological code can provide codes with more favorable properties. TN codes also provide a more general framework for better understanding holographic QEC codes such as those constructed out of isometric [167] and hyperinvariant [168] tensors that are featured in models of quantum gravity and holography, and can additionally exhibit favorable properties for practical QEC [169].
@@ V.2 Syndrome decoding
Decoding a quantum error correcting code involves processing syndrome measurements using a given theoretical error model. This process yields appropriate correction operations and involves solving the inference problem of finding the set of errors that ocurred in the computation with the highest likelihood based on the syndrome information measured and on the error model at hand. This process has been shown to be formally equivalent to contracting a TN [161, 170].
The error model of a quantum error correcting code provides two sets of information: 1) the probability of occurrence of the different physical errors and 2) the recognition of what syndrome measurements each of these errors affects and how each one affects the logical qubit(s) encoded. Given a set of syndrome measurements, there are exponentially many configurations of physical errors compatible with it. This set of error configurations can be further decomposed into several subsets, defined after their effect on the logical information encoded; these subsets are often called cosets. The sum of the joint probabilities of the error configurations of each coset is proportional to the likelihood of the coset. Optimally, a decoder infers the maximum likelihood coset as the most probable logical outcome of the code, given a set of syndrome measurements. This is an exponentially large sum of probabilities computed, in a good approximation, by the contraction of a TN.
There are two approaches to carry out this summation. The first and most popular one consists of identifying a complete set of symmetries of the coset, i.e., all sets of errors that leave the logical information and the syndrome mesasurements unchanged. In this setting, an enumeration of all error configurations included in the coset is achieved by finding a particular configuration in the set and applying to it all combinations of the symmetries [161, 170, 6, 171, 172].
A second approach consists of building a TN directly from the error model connectivity between errors and syndrome detectors. In this approach all error configurations outside a coset are not considered by explicitly zeroing out tensor entries corresponding to them. This approach has the benefit of working with more general error models directly, without the need to find a set of symmetries that is computationally beneficial [173, 174, 175].
@@ V.3 Quantum error mitigation
Error mitigation methods are expected to play a crucial role in near-term quantum experiments until fault-tolerance is fully achieved [177, 178]. One of the promising methods in this area is probabilistic error cancellation (PEC). However, it requires precise knowledge of the underlying noise model at the gate level. PEC creates an ensemble of circuits, with a denoiser, which on average replicates a noise-inverting channel, inserted according to a quasi-probability distribution, effectively canceling out the noise. A quasi-probability distribution is a probabilistic framework that assigns a sign to each sample, allowing for the representation of non-physical channels. The number of samples (measurement overhead) required to address sign cancellations can be effectively bounded by Hoeffding’s inequality, growing exponentially with the number of qubits and circuit depths. Despite the overhead, it is anticipated that error mitigation will remain practical for circuits with qubit counts in the hundreds and equivalent circuit depth [87]. The tensor-network error mitigation (TEM) algorithms represent a similar paradigm for reducing overhead. As investigated in a series of recent works [179, 180, 181, 182], TEM provides a universal lower cost bound for error mitigation, specifically a quadratic cost reduction. TEM is a post-processing procedure that acts on a set of randomized local measurements, as designed to cancel errors by incorporating classically noise-inverting channels. The mitigiated estimations are given by contracting the circuit-level TN. For systems containing more than a thousand qubits with equivalent circuit depth, we foresee the development of hybrid approaches combining quantum error correction with error mitigation techniques [183]. It remains an open question whether error mitigation alone, or in combination with quantum error correction, can lead to utility in near-term quantum computation or provide a potential practical advantage in quantum algorithms.
@@ VI Tensor networks for quantum machine learning
Over the past decade, TNs have gained significant interest in machine learning (ML). By compressing high-dimensional linear layers in neural networks, TNs reduce memory usage and the number of training parameters [184, 185, 186]. Moreover, TNs offer strong analytical interpretability and are related to various established ML techniques, such as Boltzmann machines [187, 188], Born machines [189], autoregressive models [190], and hidden Markov models [191]. The architectural flexibility of TNs allows the introduction of inductive biases into ML models based on intrinsic data topology and correlations [186, 192, 193, 194], ultimately enhancing training efficiency and model generalization. This synergy has paved the way for employing TNs across various machine-learning applications, including discriminative, generative, and regression frameworks.
@@
Although the contraction complexity of many TNs grows polynomially with the bond dimension, the degree of the polynomials is higher than three in some cases (such as for MERA), which is often regarded as inefficient in practical ML. Quantum computers may assist TN-based ML by implementing parameterized quantum circuits (PQC) reflecting specific TN architectures, resulting in TN-based quantum ML [195] (TN-QML) which avoids expensive contractions. Moreover, in the context of quantum simulation, parameterizing TNs with quantum gates has been shown to offer an exponential reduction in the number of parameters required to achieve results similar to those of classical TNs [138, 141], which could potentially translate to ML tasks as well. On the other hand, such variational TN circuits offer advantages over commonly used quantum hardware-efficient ansätze, such as being free from barren plateaus [196, 197, 145] (flat regions in the optimization landscape that impede training). Additionally, the TN structure yields qubit-efficient implementations [198] and robustness against decoherence [198, 199, 200].
@@
Consequently, quantum versions of MPS [198, 201], TTN [198, 199, 202], and MERA [199, 202] have been explored in QML for both discriminative [198] and generative [201] tasks. The TN-QML approach also inherently enables quantum applications, such as quantum phase classification directly on quantum devices using quantum convolutional neural networks (QCNNs) based on the TTN and MERA architectures [203, 202]. However, although the QCNN was initially considered a candidate for exponential quantum advantage, recent work [204] has argued against this, suggesting it can be dequantized using classical shadow techniques.
@@
However, TNs are promising in QML as a tool for encoding classical data into quantum circuits [205, 206] and for pre-training the TN-PQC on classical computers [198, 207, 208, 209] before extending to quantum ansätze that cannot be simulated classically. Finally, TN surrogate modeling can be employed to test QML moddels for dequantization [210]. For an in-depth review of using TNs in the QML, we refer the reader to [195].
@@ VII Discussion and Outlook
TNs are a valuable tool for the field of quantum computing. They can be important for analyzing computational complexity classes. For example, finding a polynomial-time tensor network algorithms for a problem [211] proves that the problems is in the complexity class $\mathcal{P}$. Similarly, one can empirically argue that a problem is hard by trying existing state-of-the-art TN techniques and observing exponential scaling. Along the same vein, TNs can help validating quantum advantage claims of finite-size quantum experiments, which may not admit asymptotic quantum advantage. Here, rigorous benchmarking of TN techniques as well as other classical techniques allows to decide whether a quantum experiment is out of reach for classical computers.
TNs have proved their value in quantum algorithm analysis, circuit synthesis, and quantum error correction and mitigation, allowing them to enhance the capabilities of quantum computation. For quantum algorithm analysis, TNs are often used to simulate a quantum algorithm and extrapolating the performance scaling. For quantum circuit synthesis, TNs allow for certain states or operators to be prepared on hardware efficiently with limited connectivity. For quantum error correction, TNs are valuable theoretical tools enabling the construction of certain families of codes. TNs can also be used for constructing optimal decoders. For quantum error mitigation, TNs enable an overhead reduction. In QML, TNs are used to design data-specific quantum models that improve resource efficiency, trainability, and robustness against noise.
TNs continue to drive progress across multiple areas of quantum computing as demonstrated throughout this review. As we transition towards fault-tolerant quantum devices, TNs will likely maintain their significance, particularly in potential hybrid quantum-classical devices which are anticipated to integrate quantum computers with specialized classical accelerators like the GPUs, TPUs, and FPGAs.
Looking forward, we foresee that TNs will evolve to include new algorithmic approaches and techniques to meet the demands of fault-tolerant quantum computers. As quantum hardware capabilities improve, the synergy between TNs and quantum computing may deepen, keeping TNs an important tool for further develoment of future quantum computing systems.

### v2

@@ IV Quantum circuit synthesis
TNs offer a promising general pathway for addressing these challenges [137]. TN architectures, such as MPS and PEPS, feature an inherently geometrical layout provided by the prescribed decomposition/representation of general tensor formats. It turns out that for circuit descriptions, which inherently are TNs, such layouts are particularly attractive due to the ease in attaining connectivity within adjacent sets of qubits. Although this feature is exploited in studying long/short range interactions in condensed matter systems, it naturally provides a partial solution to the problem of circuit synthesis in terms of realizing native hardware connectivity. By representing quantum states and operations as TNs and further casting them as circuits, one can simplify both circuit design as well as address the problem of compiling arbitrary unitaries into natively realizable gates. This can reduce the complexity of quantum circuits and enable more efficient synthesis.
@@ IV.2 Examples of a TN state and operator preparation
Using the introduced prescription, we provide examples of important TN architectures in the literature, and demonstrate how they can be mapped to quantum circuits.
Figure 3 depicts how one converts an MPS into a quantum circuit using the steps from the above subsection. The key element in this procedure is that an MPS can be exactly transformed into a canonical form, including only isometric tensors, which is a unique feature of an MPS—this is not the case for the higher-dimensional TNs like PEPS. Preparing an MPS as a quantum circuit was first introduced in [139], and has since been theoretically explored by many others [140, 141, 142, 143, 144]. An experimental realization of an MPS has been extensively explored across various quantum computer architectures [145, 146, 147, 148].
@@
The PEPS, however, generally lack a canonical exact isometric form due to the higher connectivity inherent in their underlying structure. In other words, one cannot exactly map a PEPS onto a quantum circuit without an exponential amount of postselection. However, one can create a subclass of PEPS, referred to as isoPEPS, to establish a proper connection to quantum circuits. In the isoPEPS class, each tensor, or set of tensors, respects the isometric condition, making it equivalent to a quantum circuit [91, 150, 151]. The question of understanding the variational power of isoPEPS and the way how it is compared to PEPS and to 2dMERA and how they can be realized experimentally on a digital quantum computer is now an active area of research.
Among all tensor network methods for classical simulation of quantum systems, the MERA stands out as the most natural representation of a quantum circuit: it was originally envisioned in the reverse direction of a quantum circuit in which all isometries are embedded in unitaries. It features an intrinsic robustness to noise and does not suffer from barren plateaus [152, 153, 154]. The MERA has recently been demonstrated on the ion-trapped digital quantum computer to probe the criticality of many-body systems [155, 156, 157][158].
@@ IV.3 Implementation techniques
As mentioned in Sec. 4.1, TNs lack the inherent directionality compared to real quantum circuits. Because of this, we must choose how tensors are executed in time, which then is related to the quantum circuit architecture. This freedom allows TN states to be prepared holographically. Introduced first in [145], a holographic preparation of a TN state takes a spatial dimension of the physical state and prepares it sequentially in time, with the important feature that once a circuit is done preparing one site, the qubit can be reused to prepare the next site.
@@
Beyond the exact implementations of mapping a TN to a quantum circuit described above, various variational methods are employed to prepare an approximate TN state [166, 167, 168][169, 170]. The basic idea is to numerically optimize a variational ansatz to approximate the target state, typically resulting in a shallower quantum circuit. In addition to preparing a TN state, the work [171] used variational methods to instantiate a generic quantum circuit where the TN formulation is leveraged as an efficient back-end in the compilation workflows.
@@ V Quantum error correction and mitigation
Quantum error correction (QEC) plays a critical role in safeguarding quantum information from errors caused by decoherence, dissipation, and control inaccuracies [136, 172]. In this section, we discuss two main topics: (i) the parallels between quantum error-correcting codes and TNs, along with (ii) the application of TNs in decoding, including the formal reduction of optimal decoding to a TN contraction [173].
@@ V.1 Tensor-network codes
The first major intersection between TNs and QEC was the establishment of a formal correspondence between certain QEC codes and the associated TNs [173]. Convolutional codes, concatenated block codes, and topological codes, for example, can be respectively represented by the MPS, tree TNs, and PEPS. This insight allows for importing well-established tools of TNs to represent and analyze QEC codes. In particular, the problem of an optimal decoding of an QEC code (i.e. diagnosing errors, see the discussion in Section V.2) was shown to be formally equivalent to contracting an associated TN [173]. For convolutional and concatenated block codes, which have efficiently contractable TNs, this correspondence thereby provides the efficient optimal decoders.
The QEC-TN correspondence was later generalized by the introduction of TN stabilizer codes [174, 175] and the ‘quantum LEGO’ formalism [176, 177, 178] for constructing large codes out of a finite set of smaller ‘seed’ codes. The seed codes are represented by small tensors that are combined into a TN to represent a larger code. This construction was shown to be universal in the sense that any qudit QEC code can be represented by a TN built out of three elementary seed tensors [176]. A further advantage to constructing codes in this fashion is due to the fact that TN codes come naturally equipped with optimal decoders that are evaluated by contracting a TN [174].
Quantum codes generated using tensor networks generalize code concatenation, a fact that has been observed in Refs. [176, 177]. Several examples of code constructions hailing from TN structures can be seen in topological [175] and holographic [179, 180, 181] tensor-network codes, as well as examples of non-Abelian stabilizer [182], non-additive topological [183], and approximate quantum codes [184, 185]. Apart from their usefulness in studying topologically-ordered matter or the Anti-de Sitter/ Conformal Field Theory (AdS/CFT) correspondence, many of the codes proposed exhibit favorable properties for practical QEC [186, 187, 178]. Indeed, it has been shown that machine-learning methods can even be utilized in order to search for QEC codes possessing desirable code parameters, entirely within the tensor-network formalism [188, 189].
@@ V.2 Syndrome decoding
Decoding a quantum error correcting code involves processing syndrome measurements using a given theoretical error model. This process yields appropriate correction operations and involves solving the inference problem of finding the set of errors that occurred in the computation with the highest likelihood based on the syndrome information measured and on the error model at hand. Such a decoding method is known as maximum-likelihood decoding, and has been shown to be formally equivalent to contracting a TN [173, 190], even in the case of more invasive noise models, such as those employing correlated noise [191].
The error model of a quantum error correcting code provides two sets of information: 1) the probability of occurrence of the different physical errors and 2) the recognition of what syndrome measurements each of these errors affects and how each one affects the logical qubit(s) encoded. Given a set of syndrome measurements, there are exponentially many configurations of physical errors compatible with it. This set of error configurations can be further decomposed into several subsets, defined after their effect on the logical information encoded; these subsets are often called cosets. The sum of the joint probabilities of the error configurations of each coset is proportional to the likelihood of the coset. Optimally, a decoder infers the maximum likelihood coset as the most probable logical outcome of the code, given a set of syndrome measurements, resulting in an exponentially large sum of probabilities computed in good approximation by the contraction of a TN.
There are two approaches to carry out this summation. The first and most popular one consists of identifying a complete set of symmetries of the coset, i.e., all sets of errors that leave the logical information and the syndrome measurements unchanged. In this setting, an enumeration of all error configurations included in the coset is achieved by finding a particular configuration in the set and applying to it all combinations of the symmetries [173, 190, 4, 192, 193].
A second approach consists of building a TN directly from the error model connectivity between errors and syndrome detectors. In this approach all error configurations outside a coset are not considered by explicitly zeroing out tensor entries corresponding to them. This approach has the benefit of working with more general error models directly, without the need to find a set of symmetries that is computationally beneficial [194, 195, 196]. One can also utilize tensor networks in order to enumerate error paths that appear in the realization of fault-tolerant syndrome extraction, albeit with significantly reduced computational complexity in certain specialized cases [197, 198, 177].
Both approaches to TN decoding can be successfully leveraged with diverse noise models, such as erasure [199, 200]; fractal noise [201]; depolarizing noise [202, 200]; biased noise, and explorations of quantum channel capacity via the hashing bound [202, 203, 204, 205, 206, 207]; as well as non-Markovian noise sources [208]. In addition to those noise models mentioned, circuit-level noise models [194, 209] can also be decoded, albeit approximately.
@@ V.3 Quantum error mitigation
Error mitigation methods are expected to play a crucial role in near-term quantum experiments until fault-tolerance is fully achieved [211, 212]. One of the promising methods in this area is probabilistic error cancellation (PEC). However, it requires precise knowledge of the underlying noise model at the gate level. PEC creates an ensemble of circuits, with a denoiser, which on average replicates a noise-inverting channel, inserted according to a quasi-probability distribution, effectively canceling out the noise. A quasi-probability distribution is a probabilistic framework that assigns a sign to each sample, allowing for the representation of non-physical channels. The number of samples (measurement overhead) required to address sign cancellations can be effectively bounded by Hoeffding’s inequality, growing exponentially with the number of qubits and circuit depths. Another approach of error mitigation is zero-noise extrapolation (ZNE). ZNE artificially introduces noise to obtain results at different noise levels before extrapolating to the noiseless case [211], and was experimentally demonstrated [90].
Despite the overhead, it is anticipated that error mitigation will remain practical for circuits with qubit counts in the hundreds and equivalent circuit depth [90]. The tensor-network error mitigation (TEM) algorithms represent a similar paradigm for reducing overhead. As investigated in a series of recent works [213, 214, 215, 216], TEM provides a universal lower cost bound for error mitigation, specifically a quadratic cost reduction. TEM is a post-processing procedure that acts on a set of randomized local measurements, as designed to cancel errors by incorporating classically noise-inverting channels. The mitigated estimations are given by contracting the circuit-level TN. Notably, it was experimentally demonstrated on a 91-qubit circuit with 4095 two-qubit gates. [217] For systems containing more than a thousand qubits with equivalent circuit depth, we foresee the development of hybrid approaches combining quantum error correction with error mitigation techniques [218]. It remains an open question whether error mitigation alone, or in combination with quantum error correction, can lead to utility in near-term quantum computation or provide a potential practical advantage in quantum algorithms.
@@ VI Tensor networks for quantum machine learning
Over the past decade, TNs have gained significant interest in machine learning (ML). By compressing high-dimensional linear layers in neural networks, TNs reduce memory usage and the number of training parameters [219, 220, 221]. Moreover, TNs offer strong analytical interpretability and are related to various established ML techniques [222, 223, 224, 225, 226]. The structural flexibility of TNs allows the introduction of inductive biases (constraints on the learning algorithm to restrict the space of possible models to plausible ones) into ML models based on intrinsic data topology and correlations [221, 227, 228, 229], ultimately enhancing training efficiency and model generalization to data unseen during training. This synergy has paved the way for employing TNs across various ML applications, also sometimes referred to as quantum-inspired ML.
Although the contraction complexity of many TNs grows polynomially with the bond dimension, the degree of the polynomials is higher than three in some cases (such as for MERA), which is often regarded as inefficient in practical ML. Quantum computers can potentially address this problem of TN-based ML by implementing parameterized quantum circuit (PQC) ansätze reflecting specific TN architectures such as MPS [230, 231], TTN [230, 232, 233], and MERA [232, 233], resulting in TN-based quantum ML [234] (TN-QML).
In regression and classification tasks [230], TN-QML obtains the output of a TN neural layer by measuring observables on the corresponding quantum state rather than performing costly tensor contractions on classical hardware. In turn, in generative tasks [231], TN-QML generates data by directly sampling from the quantum state encoding the TN, thereby avoiding expensive classical sampling that would otherwise again rely on tensor contractions.
@@
Moreover, in the context of quantum simulation, parameterizing TNs with quantum gates has been shown to offer an exponential reduction in the number of parameters required to achieve results similar to those of classical TNs [146, 149], which could potentially translate to ML tasks as well. On the other hand, such variational TN circuits offer advantages over commonly used quantum hardware-efficient ansätze, such as being free from barren plateaus [235, 236, 153] (flat regions in the optimization landscape that impede training). Additionally, the TN structure yields qubit-efficient implementations [230] and robustness against decoherence [230, 232, 237].
@@
The TN-QML approach also inherently enables quantum applications, such as quantum phase classification directly on quantum devices using quantum convolutional neural networks (QCNNs) based on the TTN and MERA architectures [238, 233]. However, although the QCNN was initially considered a candidate for exponential quantum advantage, recent work [239] has argued against this, suggesting it can be dequantized using classical shadow techniques.
@@
Additionally, TNs are promising in QML as a tool for encoding classical data into quantum circuits [240, 241] and for pre-training the TN-PQC on classical computers [230, 242, 243, 244] before extending to quantum ansätze that cannot be simulated classically. Finally, TN surrogate modeling can be employed to test QML models for dequantization [245]. For an in-depth review of using TNs in the QML, we refer the reader to [234].
@@ VII Discussion and Outlook
TNs are a valuable tool for the field of quantum computing. Moreover, they can be important for analyzing computational complexity classes. For example, finding a polynomial-time tensor-network algorithm for a problem [246] proves that the problem is in the complexity class $\mathcal{P}$. Similarly, one can empirically argue that a problem is hard by trying existing state-of-the-art TN techniques and observing exponential scaling. Along the same vein, TNs can help validating quantum advantage claims of finite-size quantum experiments, which may not admit asymptotic quantum advantage. Here, rigorous benchmarking of TN techniques as well as other classical techniques allows to decide whether a quantum experiment is out of reach for classical computers.
In particular, recent developments in TNs and other classical simulation techniques have significantly improved our ability to simulate quantum many-body systems [247], as well as near-Clifford circuits. These methods arise from representing TN states directly in the Pauli basis [248]. Recent advancements in representing TN states in both a Pauli basis as well as a computational basis have shown promising results [249], and it remains an open and interesting question on what classes of circuits can be classically simulated exactly or approximately. Additionally, there is potential in understanding the quantum circuit synthesis of TN states given that they are in the Pauli basis.
On quantum error correction, one future direction may be utilizing the quantum LEGO formalism to write down families of quantum low-density parity-check codes [250] in a systematic fashion. This is an important direction since quantum low-density parity-check codes represent a promising avenue in quantum error correction [251]. Additionally, the first explorations of non-Abelian stabilizer and non-additive codes have already been constructed [183, 182]; it would be interesting to see whether or not the quantum LEGO formalism or further extensions can provide insight into systematic constructions of such codes. Moreover, highly accurate TN decoding is too slow for real-time hardware decoding [210]; as such, it could be valuable to design heuristic algorithms that can form approximate TN decoding well. The alternative could be using AI for decoding [252, 253], which may benefit from noisy simulations using tensor networks for model training.
Finally, using TNs in QML requires both further investigation into their use in classical ML and the development of new advances in QML itself. Although recent studies question the feasibility of achieving a practical exponential quantum advantage with QML [254, 255], it remains to be seen whether TN-QML can offer any polynomial advantage in this context.
TNs continue to drive progress across multiple areas of quantum computing as demonstrated throughout this review. As we transition towards fault-tolerant quantum devices, TNs will likely maintain their significance, particularly in potential hybrid quantum-classical devices which are anticipated to integrate quantum computers with specialized classical accelerators like the GPUs, TPUs, and FPGAs. Looking forward, we foresee that TNs will evolve to include new algorithmic approaches, integration with AI, and techniques to meet the demands of fault-tolerant quantum computers. As quantum hardware capabilities improve, the synergy between TNs and quantum computing may deepen, keeping TNs an important tool for further development of future quantum computing systems.